Rs.16,000 is given to a person for 2 years on 10% interest what be the compound interest
Answer by stanbon(75887)

Rs.16,000 is given to a person for 2 years on 10% interest what be the compound interest
---------------
A(t) = P(1+(r/n))^(nt)
----
A(2) = 16000(1+(0.1/1))^(1*2)
----
A(2) = 16000(1+0.1)^2
--
A(2) = 16000(1.1^2)
A(2) = 19360
---
Interest earned = 19360-16000 = 3360
======================================
Cheers,
Stan H.